A hot air ballon (with bucket) has a mass of 900 kg. The balloon has a buoyant force upward of 12,000 N and a westward wind force of 500 N. Find the acceleration of the balloon. What is the force of gravity (weight) of the balloon and bucket?

show work and answer
1
Expert's answer
2010-11-01T04:28:20-0400
The force of gravity of the baloon with bucket is F = mg = 900 *10 = 9,000 N - downward.
Let's find the vector sum of all forces:
√ (12,000-9,000)2 + 5002 = 3041 N.
a = F/m = 3041[N]/900[kg] = 3.4 [m/sec2]